OpenStudy (anonymous):

how do i evaluate 1 ) the integral of 3^X dx 2)the integral from -2 to 0 of (3^3-5^2)

OpenStudy (mimi_x3):

\[\large \int\limits3^{x} dx \]

OpenStudy (mimi_x3):

You have to like change the base so... \[\large e^{\ln3*x} \]

OpenStudy (anonymous):

oh i c , i really need a good website that has all different types of problems our book is horrible

OpenStudy (mimi_x3):

wait..i made a typo.. \[\large \frac{1}{\ln3} *3^x + C=>\frac{3^{x}}{\ln3} +C\]

OpenStudy (anonymous):

im in calc how do you remember and know which formats to use?

OpenStudy (mimi_x3):

Hm..I used the formula.. change of the base. \[\large a ^{bx + c} = e ^{lna*(bx+c)}\] Primitive: \[\large \frac{1}{blna} *a^{bx+c} \] It's kinda confusing..probably you have a better one. xD

OpenStudy (anonymous):

ya this is the one our book has thanks i just have trouble ordering it when they get complicated

OpenStudy (mimi_x3):

ordering it..?

OpenStudy (anonymous):

do you know a website that has this info in simple explenations? i dont get the book descriptions at all

OpenStudy (anonymous):

i can copy similar problems but i dont rlly get what its meaning so i can work more complex ones

OpenStudy (mimi_x3):

No sorry; those formulas are from my book. What don't you get?

OpenStudy (anonymous):

how would u go about the second example?

OpenStudy (mimi_x3):

I have to go now; sorry. Try posting it again, someone might be able to help you.

OpenStudy (anonymous):

alright thanks